Mass of KxFe(C2O4)y · zH2O : 4.70 g

Mass of sample : 0.175 g

Mass of FeCl3 used in preparation : 1.60 g

Molarity of standard NaOH used : 0.100

V1, volume of standard NaOH required for first equivalence point : 7.750 mL

V2, volume of standard NaOH required for second equivalence point : 19.70 mL

1. Calculate the mass of potassium in the sample

2. Calculate the percent of potassium in the sample

3. Calculate the volume of standard NaOH that reacted with iron

4. Calculate the mass of iron in the sample

5. Calculate the percent of iron in the sample
Answer #1

Mass of KxFe(C2O4)y · zH2O : 4.70 g

Mass of sample : 0.175 g

Mass of FeCl3 used in preparation : 1.60 g

Molarity of standard NaOH used : 0.100

V1, volume of standard NaOH required for first equivalence point : 7.750 mL

V2, volume of standard NaOH required for second equivalence point : 19.70 mL

Answer these questions using info above:

Calculate the mass of potassium in the sample :

1st equivalence point corresponds to K+ ;

mol K+ = moles of NaOH used in titration (from V1 ) = 0.100 mol/L * 7.75 ml/1000 ml/L = 0.000775 moles

mass of K+ in sample = 0.000775 moles* 39.1 g/mol = 0.0303 g

Calculate the percent of potassium in the sample :

mass of K+ in sample / mass of sample *100

0.03030 g / 0.195 g *100 = 17.31 %

Calculate the volume of standard NaOH that reacted with iron :

V2 -V1 = 19.70 ml - 7.750 ml =11.95 mL

Calculate the mass of iron in the sample :

three moles of hydroxide ion are required to react with each mole of iron ion in the salt.

moles of Fe = 1/3* moles of NaOH used in titration (form V2 - V1) = 0.100 mol/L * 7.75 ml/1000 ml/L /3 = 0.000256 moles
/mass of Fe in sample = (moles of Fe)(55.85 grams/mole) = 0.000256 moles*55.85 g/mol = 0.0142 g

Calculate the percent of iron in the sample :

mass of Fe in sample / mass of sample *100

0.0163 g / 0.195 g *100 =7.28 %